OpenStudy (anonymous):

The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are green? Color Number of Pulls red 6 green 4 orange 7 black 3

OpenStudy (kropot72):

Total number of pulls = 20 Number of green pulls = 4 Experimental proportion of green = 4/20 Estimated number of green marbles in bag = 50 * (4/20) = ?